OpenStudy (anonymous):

An object is blastedupward at initial velocity of 240 ft/s the h(t)=vot-16t2 how long will it take the object to hit the ground after takeoff

OpenStudy (amistre64):

-16t^2 +Vot +So ; So=0 since it starts at the ground the roots to this are: t(-16t +240); t=0 and t=240/16 right?

OpenStudy (amistre64):

15 sec

OpenStudy (anonymous):

Preciate it

OpenStudy (amistre64):

your welcome :)